OpenStudy (lolofranky):

I need help on this question: The real root of the equation x3 – x + 4 = 0 is –1.796 to three decimal places. Determine the real root for each of the following. (a) (x – 1)3 – (x – 1) + 4 = 0 (b) 8x3 – 2x + 4 = 0

OpenStudy (lolofranky):

the three on the side is meant to be cubed

ganeshie8 (ganeshie8):

Hint for part a : ` f(x-k) ` shifts the graph of ` f(x) ` to right by `k` units.

ganeshie8 (ganeshie8):

Hint for part b : ` f(k*x) ` shrinks the graph of ` f(x) ` horizontally by a factor of `k` when k > 1.

OpenStudy (mathmate):

Further hint: If f(x)=x^3-x+4 8x^3-2x+4=(2x)^3-(2x)+4=f(2x)

OpenStudy (lolofranky):

thank you